killerjay556:

A cylinder has a base diameter of 4in and a height of 7in. What is its volume in cubic in, to the nearest tenths place?

darkknight:

\[V= \pi*r^2*h\] \[2r=d\] h= height d = diameter r = radius V= volume

m4nt1c0r3:

No, that wouldn't be the right answer. Using the equation, \[V = \Pi \times r^2 \times height\] and \[d = 2r\] You would find that the radius is 2 and implementing that into the equation for Volume, you would get \[V = \pi \times 2^2 \times7\] and once solved, it would be \[28\pi\] or 88 cubic inches.